A pulley with a radius of 9 inches is to be used to lift a box from the ground to a platform. The rope being used has an entire length of 42 feet. If the box needs to be raised 12 feet, how many revolutions will the pulley make? How many radians is that?
Answer by venugopalramana(3286)

A pulley with a radius of 9 inches is to be used to lift a box from the ground to a platform. The rope being used has an entire length of 42 feet. If the box needs to be raised 12 feet, how many revolutions will the pulley make? How many radians is that?
PER REVOLUTION THE ROPE TTRAVELS BY ONE CIRCUMFERENCE OF THE PULLEY=2PI*9/12 FEET
IT HAS TO RAISE BY 12 FEET..
HENCE NUMBER OF REVOLUTIONS NEEDED = 12/(2PI*9/12)=12*12/18PI=8/PI=2.55 OR NEARLY 3 REVOLUTIONS.
IT AMOUNTS TO (8/PI)*2PI=16 RADIANS